# REvolve: Reward Evolution with Large Language Models for Autonomous Driving

### v1

@@ 5 Experiments
For our experiments, we used the AirSim simulator [48], a high-fidelity simulator for AD research. We used the GPT-4 Turbo (1106-preview) model as the reward designer $G$, and provided it with (a subset of) observation and action variables originating from the AirSim simulator [^2] To prevent potential plagiarism through context memory [5], we use unique environment variable names and deliberately avoid mentioning the AirSim simulator in inputs to GPT-4., as illustrated in Figure 1.
@@ 5.1 Baselines
Human Driving. To benchmark the learned policies against human driving, we recruited participants to drive in a highly realistic setup, conjoining the AirSim (simulated) environment with a Logitech driving simulator. We collected data from 3 participants, each with 10 test runs. Participants were allowed trial runs to familiarize themselves with the setup.
Expert Designed. We used an expert reward function based on best practices established in the AD literature [44, 52]. The total reward is calculated through a weighted aggregation of all the components, with the weights optimized via an exhaustive grid search conducted within the AirSim simulator in order to find the reward function that produces the best policy. Details on the expert-designed reward function can be found in Appendix C.
Eureka. Closest to our method is Eureka [30], which creates a population of reward functions per iteration and greedily iterates on the best individual based on feedback. We employed the human fitness function to refine the best individual with natural language feedback, akin to REvolve.
Additionally, to emphasize the importance of human fitness function, in general, and human feedback, in particular, we explored an automatic feedback mechanism originally proposed in Eureka. Here, the (string) feedback is composed of the statistics of all reward components tracked from intermediate training checkpoints. e.g., "collision_penalty: [-1, -0.8, -0.4, …]". Ideally, a component is deemed beneficial if its statistics consistently grow/fall during training and potentially less useful if they are constant. Moreover, the fitness score is computed using a manually designed fitness function, which evaluates the policy on key performance indicators like collisions (Boolean), speed (real), and distance covered (real). We refer to Appendix B.2 for further details. We call these variants of Eureka and REvolve, Eureka Auto and REvolve Auto, respectively. Lastly, we also included T2R [61] in our comparison, which outputs a single reward function and iteratively refines it using feedback (T2R Auto).
Figure 4: Figure compares REvolve with baselines using manually designed fitness scores. It can be observed that REvolve consistently improves over generations, achieving its highest fitness score $\sigma=0.84$. The upper limit is marked by Human Driving at $\sigma=0.90$ while the Expert Designed fitness score is fixed at $\sigma=0.65$.
Figure 5: Figure compares policy training with REvolve and Expert Designed reward functions, on episodic steps. Policy training converges at a higher value for REvolve signifying a higher success rate. This is true for REvolve and REvolve Auto, with the former outperforming the latter.
5.2 Training Details
The observation of a policy includes both visual (egocentric camera view) and sensor data (e.g. yaw, pitch, linear velocity). The action space $A$ consists of steering and throttling controls, with a size $|A|=66$. The objective is to navigate an urban environment for 1000 consecutive time steps. A successful termination of the episode happens when the agent meets the objective. Failures like collisions or breaking speed thresholds will also terminate the episode. The specific details of the RL setup is outlined in Appendix A.
REvolve Hyperparameters: We set the number of generations to $N=5$ and individuals per generation to $K=16$, similar to Eureka. The number of sub-populations was set to $I=13$, and the mutation probability was set to $p_{m}=0.5$. For each generation, $16$ policies were trained (corresponding to each reward function) over $5\times 10^{5}$ training steps, on 16 NVIDIA A100 GPU (40GB). The training time for each policy was $\approx 50$ hours. Due to the extensive computational demands and significant training times, we report on a single random seed for both REvolve and Eureka baselines. Regarding human feedback, we engaged 10 human evaluators per generation, each assessing 20 pairs of videos. No personal data was collected.
@@ 5.3 Experimental Questions & Results
How does REvolve fair against the baselines on manually designed fitness scores?
@@
As shown in Figure 4, REvolve exhibits continuous improvement across successive generations, ultimately achieving a higher fitness score on the manually designed fitness scale. Furthermore, both REvolve ($\sigma=0.84$) and REvolve Auto ($\sigma=0.81$) outperform their Eureka counterparts ($\sigma=0.72$ and $\sigma=0.70$, respectively), indicating that evolutionary search strategies yield better results than greedy search methods [^3] Note that these results are not displayed on the human fitness scale because the Elo ranking system evaluates individuals relative to one another. Therefore, comparisons between individuals from different baselines are only feasible if we directly pair them against one another.. We can also see that all methods but T2R Auto outperform the Expert Designed reward function ($\sigma=0.65$). Furthermore, we also observe the benefit of providing human feedback (REvolve: $\sigma=0.84$, Eureka: $\sigma=0.72$), as both REvolve and Eureka with provided human feedback outperforms their automatic feedback counterparts (REvolve Auto: $\sigma=0.81$, Eureka Auto: $\sigma=0.70$). None of the methods, however, surpassed the Human Driving baseline with $\sigma=0.90$, calculated by averaging the fitness scores over 30 runs (from 3 participants).
Do human fitness scores induce human-aligned behavior?
We further compared REvolve and Eureka on human evaluations. To this end, we generated rollouts from the best performing policies ($\pi^{\ast}$) from REvolve and Eureka, along with Human Driving. We paired them against each other and used the human feedback interface (as seen in Figure 1) for collecting human preference data. Collected human preference data was subsequently passed through our Elo rating system to rank each behavior. As shown in Table 1, REvolve policies secure second and third ranks, surpassed only by Human Driving.
We show the best (i.e. fittest) REvolve-designed reward functions $R^{\ast}$ in Appendix F.
Do REvolve reward functions lead to more episodic steps?
Episodic steps denote the number of actions taken by an agent within an episode before termination– either by reaching the episode length ($10^{3}$ steps) or by taking a non-recoverable action (e.g. crashing). In Figure 5 we see that REvolve-designed rewards converge to a higher number of episodic steps compared to expert-designed rewards, signifying a higher rate of successful actions per episode.
Aside from our standard environment (seen in Figure 6), we assessed the generalizability of the best REvolve-designed reward functions, $R^{\ast}$ in two new AirSim environments. Specifically, we trained policies from scratch using $R^{\ast}$ in two new scenarios: (Env 1) featuring lanes and a completely altered landscape while maintaining similar traffic conditions as our standard environment, and (Env 2) characterized by increased traffic with multiple cars actively maneuvering. From Table 1, we observe that in both environments, REvolve outperforms expert-designed rewards. However, the generalizability was lower for Env 2. This issue may stem from the temporal processing limitations of the RL policy’s visual module, which sometimes fails to detect oncoming cars. Implementing more advanced models could enhance detection capabilities [18]. Figure 7 in the Appendix A provides visual representations of these environments.
@@
Table 1: The left table shows the best policies $\pi^{\ast}$ from each framework ranked based on human preferences, alongside Human Driving. We use the Elo rating system to map preferences to scores. All policies are initially assigned an Elo score of 1500. The right table shows the generalization scores in different environments. Here Fitness denotes manually designed fitness scores (mean ± std). Env 1 and Env 2 have 1500 and 1000 episodic steps, respectively.
Rank Baseline Elo Score
1 Human Driving 1595
2 REvolve 1534
3 REvolve Auto 1518
4 Eureka 1453
5 Eureka Auto 1401
@@
5.4 Limitations.
As stated in § 5.2, due to its (deep) RL nature, training a policy given a reward function is rather computationally expensive and time-consuming. While the computational cost of REvolve is the same as for Eureka – with no resources wasted on discarding individuals – REvolve does incur a slightly higher storage cost due to the maintenance of the reward database. Another limitation of REvolve is its reliance on the closed source GPT-4 model. This was a necessary choice as existing works have shown that a considerable gap exists in RD abilities between GPT-4 and open-source LLMs [30, 61]. With rapid advancements made by the open-source community, we believe this gap will be bridged in the near future.
@@ 6 Conclusion & Future Work
We formulated RDP as a search problem using evolutionary algorithms for searching through the space of reward functions (§ 2). With REvolve, we proposed a novel evolutionary framework that integrates LLMs as intelligent operators refining solutions through natural language feedback (§ 3). While we demonstrate our framework on the challenging problem of AD (§ 5), we also note that the generality of the formulation allows for broad applicability, especially in better alignment of deep learning models with human values and objectives. Unfortunately, the proposed framework can also be exploited by bad actors.
@@
A key future question for REvolve is its real-world applicability. One promising approach involves integrating REvolve with Sim2Real algorithms [1, 20] by first learning a policy in simulation, using REvolve-designed rewards, and then transferring it to the real world. Recently, Ma et al. [31] have demonstrated that such an approach is viable for Eureka by tackling a quadruped balancing task on a physical robot. Moreover, our framework opens up opportunities for future research in Human-AI interaction. In particular, exploring the impact of more open-ended feedback from various perspectives, such as egocentric or third-person views.
Acknowledgement
This work was fully supported by the Knut and Alice Wallenberg foundation via the Wallenberg AI, Autonomous Systems and Software Program (WASP) and the Wallenberg scholar grant. We also acknowledge the Berzelius cluster at the National Supercomputer Centre, Sweden, for granting access to high-performance computing resources under projects "Berzelius-2024-178" and "Berzelius-2023-348". Finally, we thank Luc De Raedt for his valuable feedback in refining the manuscript.
Appendix for "REvolve: Reward Evolution with
Large Language Models for Autonomous Driving"
@@ Acknowledgement
The appendix is organized as follows: 1. RL Preliminaries § A where we elaborate on the reinforcement learning framework. 2. Fitness Function § B where we outline the fitness score computation from human feedback. Additionally, we outline the automatic feedback approach from Eureka. 3. Expert Designed Reward Function § C which provide details of the expert designed reward and grid search parameters. 4. Prompts § D: System and operator (mutation, crossover) prompts for GPT-4 for REvolve and REvolve Auto 5. Human Feedback Examples § E to show how human feedback allows GPT-4 to refine the reward function individual. 6. Additional Results § F provides best reward functions generated from REvolve, and some additional results.
@@
We also include our code and demonstrations of policies trained on best REvolve-designed reward functions, in the supplementary material.
Appendix A RL Preliminaries
A.1 Setup
@@
Simulation Environment. The experiments are conducted in the neighborhood environment, a high-fidelity simulator developed for autonomous vehicle research [48]. This urban simulation includes static features such as parked cars and dynamic obstacles like cars and moving animals, enabling a thorough evaluation of the RL agent in varied scenarios.
@@
Figure 6: AirSim simulator environment used for our AD experiments. The egocentric camera view displayed on the bottom left serves as the observations for the AD agent.
@@
Figure 7: Airsim generalization environments used for assessing the generalizability of REvolve-designed reward functions. The left and right pictures visualize Env 1 and Env 2 respectively.
A.2 Algorithm
We adopt the Clipped Double Deep Q-Learning approach [12], a refined version of Double Q-Learning [56]. This method uses two separate neural networks to reduce overestimation bias by calculating the minimum of the two estimated Q-values, thus providing more reliable and stable training outcomes.
In addition to Clipped Double Q-Learning, our framework integrates the Dueling Architecture [60], which improves the estimation of state values by decomposing the Q-function into two streams: one that estimates the state value and another that calculates the advantages of each action. Such structure allows the model to learn which states are valuable without the need to learn each action effect, for each state.
Furthermore, we utilize Prioritized Experience Replay (PER) [47] to focus on significant learning experiences. PER prioritizes transitions with higher Temporal Difference errors for replay, which accelerates learning and increases the efficiency of the memory used during training. Table 2 lists the RL model hyperparameters. Finally, inspired by foundational research, we have adapted our architecture to incorporate vision-based methods and the Dueling Architecture from [33] and [60], enhancing temporal dynamics and action-value precision.
Table 2: DDQN Hyperparameters
@@ A.2 Algorithm
Appendix B Fitness Function
In this section, we provide details of (1) REvolve fitness function which is based on human evaluations; (2) the automatic feedback approach proposed in Eureka.

### v2

@@ 5 Experiments
5.1 Experimental Setup
We performed experiments across three different environments: autonomous driving, humanoid locomotion, and adroit hand manipulation. For autonomous driving, we used the high-fidelity AirSim simulator (Shah et al., 2018). For humanoid locomotion (Tassa et al., 2012) and adroit hand manipulation (Rajeswaran et al., 2017), we used the MuJoCo simulator (Todorov et al., 2012). As a reward designer $G$, we used GPT-4 Turbo (1106-preview) model, which was provided (a subset of) observation and action variables originating from the simulators [^2] To prevent potential plagiarism through context memory (Carlini et al., 2021), we use unique environment variable names and deliberately avoid mentioning the simulator in inputs to GPT-4., as illustrated in Figure 1. The reward designer is prompted to provide a reward taking the functional form $r=\sum_{i}r_{i}$. We call the $r_{i}$’s reward components and the reward designer decides on their individual functional form, as well as the number of reward components. This follows the protocol established by Ma et al. (2024a).
5.2 Training Setup
Table 1: Summary of observation and action spaces, along with objectives and termination criteria for the different RL training tasks.
Task Observation Space Action Space Objective & Termination
Autonomous Driving $\mathbb{R}^{307206}$ discrete, $|A|=66$ Navigate for 1000 steps
Humanoid Locomotion $\mathbb{R}^{376}$ continuous, $A\in\mathbb{R}^{17}$ Run upright for 1000 steps
Adroit Hand Manipulation $\mathbb{R}^{39}$ continuous, $A\in\mathbb{R}^{30}$ Open the door in 400 steps
For the evolutionary search, we set the number of generations to $N=5$ and individuals per generation to $K=16$. This mimics the training setup of (Ma et al., 2024a). The number of sub-populations was set to $I=13$, and the mutation probability was set to $p_{m}=0.5$.
For each generation, $16$ policies were trained (one per reward function). In the AirSim environment, we trained Clipped Double Deep Q-Learning  (Fujimoto et al., 2018) agents for $5\times 10^{5}$ training steps per generation per agent. For the MuJoCo environments, we trained Soft Actor-Critic (Haarnoja et al., 2018) agents with $5\times 10^{6}$ steps per generation per agent. We summarize the different RL tasks in Table 1 and provide detailed descriptions of the environments and RL setups in Appendix A.
Parallel training of a single generation on 16 NVIDIA A100 GPUs (40GB) consumed approximately 50 hours for the AirSim environment and 24 hours for the MuJoCo environments. Due to these extensive computational demands, we report results using two random seeds.
In addition to REvolve (with human feedback), we also include a version in the experimental evaluation that we dub REvolve Auto. Here, the human feedback is replaced with automatically generated feedback comprised of statistics on the reward components tracked during RL training. Moreover, the fitness score, evaluating the individual agents, is computed using a manually designed fitness function. This fitness function measures an agent’s successful completion of the task at hand. For instance, the fitness score for the humanoid locomotion task encodes that the agent should not fall over and that it should attain a high average velocity. We formally describe the fitness scores for the three different environments in Appendix B.2. This automated feedback mechanism is adopted from (Ma et al., 2024a).
@@ 5.3 Baselines
Expert Designed. For autonomous driving, we used an expert reward function based on best practices established in the literature (Spryn et al., 2018). The total reward is calculated through a weighted aggregation of all the components, with the weights optimized via an exhaustive grid search conducted within the environments to find the reward function that produces the best policy. For humanoid locomotion and adroit hand manipulation, we used expert rewards provided by Gymnasium and Gymnasium-Robotics [^3] https://gymnasium.farama.org,https://robotics.farama.org, respectively. Details on the expert-designed reward functions can be found in Appendix C.
Eureka. We also compare REvolve to Eureka, cf. Section 4. Akin to REvolve, we employed the human fitness function to refine the best individual with natural language feedback. Henceforth, we denote it as Eureka. We also include the original version of Eureka with automated feedback, which we denote by Eureka Auto. Training times are identical to the ones for REvolve.
Figure 3: We plot the fitness score – over 2 random seeds – of the best-performing policy per generation for autonomous driving (left), humanoid locomotion (middle), and adroit hand manipulation (right), cf. Table 1. Each plot also contains the fitness score obtained using an expert-designed reward function. The fitness score for the human driving in the autonomous driving setting (left) was obtained by observing the driving behavior of an expert human conducting driving sessions in a combined setting of AirSim and the high-fidelity Logitech driving simulator. We treat this as the upper limit of performance. We give the formal definitions of the fitness functions for each of the three environments in Equations 2, 3, and 4, respectively (in Appendix B.2).
@@ 5.4 Experimental Questions & Results
As shown in Figure 3, REvolve exhibits continuous improvement over successive generations, ultimately achieving a higher fitness score on the manually designed fitness scale. Furthermore, both REvolve and REvolve Auto outperform their Eureka counterparts, indicating that evolutionary search strategies yield better results than greedy search methods [^4] Note that these results are not displayed on the human fitness scale because the Elo ranking system evaluates individuals relative to one another. Therefore, comparisons between individuals from different baselines are only feasible if we directly pair them against one another.. We also show that all methods outperform the Expert Designed reward function. Furthermore, we observe the benefit of providing human feedback, as both REvolve and Eureka with provided human feedback outperform their automatic feedback counterparts.
How does REvolve fair against the baselines?
We compare REvolve to the baselines using the task-specific fitness score described in Appendix B.2 and report the results in Figure 3. We observe that REvolve (with human feedback) consistently outperforms the other methods across all three environments and is only beaten by the expert human driver in Autonomous Driving, cf. Figure 3 (left). For the interested reader, we give the best (i.e., fittest) REvolve-designed reward functions $R^{\ast}$ in Appendix G.
We also note that in the Humanoid Locomotion and Adroit Hand tasks REvolve Auto (i.e., with automated feedback) performs on par with Eureka (with human feedback), despite the former’s evolutionary setting. This could be attributed to the limited and sparse nature of automated feedback (velocity for humanoid, steps to succeed for adroit hand) in contrast to the more detailed human feedback in Eureka. In contrast, automated feedback in Autonomous Driving is dense and incorporates multiple factors like collision, speed, and distance (Appendix B.2) – leading to bigger improvements during the evolutionary search. This also highlights that the combination of evolutionary search and human feedback is crucial for optimizing reward functions.
@@
Table 2: The table shows the best policies $\pi^{\ast}$ from each framework ranked based on human preferences for each environment. Specifically, we use rollouts from the best policies of each framework and gather human preferences on the paired rollouts, as described in Section 3.3. Then, we use the Elo rating system to map preferences to scores. All policies are initially assigned an Elo score of 1500.
Environment Rank Baseline Elo Score
Autonomous Driving 1 Human Driving 1586
2 REvolve 1575
3 REvolve Auto 1529
4 Eureka 1499
5 Eureka Auto 1411
6 Expert Designed 1397
Humanoid Locomotion 1 REvolve 1586
2 Eureka 1557
3 REvolve Auto 1514
4 Eureka Auto 1434
5 Expert Designed 1407
Adroit Hand Manipulation 1 REvolve 1594
2 REvolve Auto 1524
3 Eureka 1522
4 Eureka Auto 1443
5 Expert Designed 1415
How do humans judge REvolve policies?
We further compared REvolve and Eureka on human evaluations. To this end, we generated roll-outs from the best performing policies ($\pi^{\ast}$) from REvolve and Eureka. We paired them against each other and used the user feedback interface (Figure 1) for collecting preference data. This data was subsequently passed through an Elo rating system to rank each behavior, cf. Appendix B.1. As shown in Table 2, REvolve policies again outperform Eureka policies across all three tasks, and it is only the human expert driver beating REvolve in Autonomous Driving.
We have submitted example rollout videos as part of the supplementary material.
5.5 Ablation Studies
Figure 4: Comparison of REvolve’s performance on Autonomous Driving using mutation only, crossover only, and both combined.
Table 3: Table shows the generalization scores in unseen AirSim environments. Fitness denotes manually designed fitness scores (mean ± std). Env 1 & Env 2 have 1500 & 1000 episodic steps, respectively.
@@ 5.5 Ablation Studies
Eureka 0.57 ± 0.05 99 ± 28
How do different genetic operators impact overall performance?
We evaluated REvolve’s performance in Autonomous Driving by running evolution search with three setups: using only mutation, only crossover, and the combination of both. As shown in Figure 4, REvolve with mutation & crossover outperforms REvolve with only crossover, which in turn outperforms REvolve with just mutation.
Aside from our standard Autonomous Driving environment (seen in Figure 5), we assessed the generalizability of the best REvolve-designed reward functions $R^{\ast}$ in two new (AirSim) environments. Specifically, we trained policies from scratch using $R^{\ast}$ in two new scenarios: (Env 1) featuring lanes and a completely altered landscape while maintaining similar traffic conditions as our standard environment, and (Env 2) characterized by increased traffic with multiple cars actively maneuvering. From Table 3, we observe that in both environments, REvolve outperforms Eureka and expert-designed rewards. However, the generalizability was lower for Env 2. This issue may stem from the temporal processing limitations of the RL policy’s visual module, which sometimes fails to detect oncoming cars. Implementing more advanced models could enhance detection capabilities Hazra et al. (2023). Figure 6 provides visual representations of these environments, while Figure 13 provides episodic steps on the Y-axis over total steps on the X-axis during training for both generalization environments. We did not conduct similar experiments on the MuJoCo simulator, as the Humanoid and Adroit Hand environments are limited to only a single variant.
5.6 Limitations
As stated in Section 5.2, due to its (deep) RL nature, training multiple policies is rather computationally expensive and time-consuming. That said, the computational cost of REvolve is the same as for Eureka – with no resources wasted on discarding individuals – REvolve does incur a slightly higher storage cost due to the maintenance of the reward database (see Appendix D). Another limitation of REvolve is its reliance on the closed source GPT-4 model. This was a necessary choice as existing works have shown that a considerable gap exists in reward design abilities between GPT-4 and open-source LLMs (Xie et al., 2024; Ma et al., 2024a). With rapid advancements made by the open-source community, we believe this gap will be bridged in the near future.
@@ 6 Conclusion & Future Work
We formulated the RDP as a search problem using evolutionary algorithms for searching through the space of reward functions (Section 2). With REvolve, we proposed a novel evolutionary framework that leverages human preferences as fitness scores to guide the search process and integrates LLMs as intelligent operators to refine outputs through natural language feedback (Section 3). Our results show that this evolutionary search-based framework considerably outperforms the state of the art across three different simulation environments.
@@
A key future question for REvolve is its real-world applicability. One promising approach involves integrating REvolve with Sim2Real algorithms Allamaa et al. (2022); Hu et al. (2024) by first learning a policy in simulation, using REvolve-designed rewards, and then transferring it to the real world. Recently, Ma et al. (2024b) have demonstrated that such an approach is viable for Eureka by tackling a quadruped balancing task on a physical robot. Moreover, our framework opens up opportunities for future research in Human-AI interaction. In particular, exploring the impact of more open-ended feedback from various perspectives, such as egocentric or third-person views.
Ethics statement
During human feedback experiments, we engaged 10 human evaluators per generation, each assessing 20 pairs of videos. Participants were recruited on a volunteer basis and the preference learning task was clearly explained to them. We note that no personal data was collected, and participants were compensated fairly per institutional guidelines. Our study adhered to ethical standards, followed institutional guidelines, and did not involve any potential risks related to privacy, security, or discrimination.
Reproducibility
In the Appendix, we provide extensive details on our experimental protocol. Specific details can be found in the respective sub-appendices as detailed below. The reinforcement learning framework, including the algorithm, inputs, outputs, and hyperparameters, is outlined in Appendix A. The fitness functions are thoroughly discussed in Appendix B. The expert-designed rewards are explained in Appendix C. Detailed prompt examples for GPT-4 can be found in Appendix E. Finally, we list the best reward functions generated from REvolve in Appendix G. We will open-source our code shortly.
Appendix
@@
The appendix is organized as follows: Section A Reinforcement Learning Preliminaries, where we elaborate on the reinforcement learning framework for all environments; Section B Fitness Function, where we outline the fitness score computation from human feedback, as well as outline Eureka’s automatic feedback approach; Section C Expert Designed Reward Function, which provides details of the expert-designed reward and grid search parameters; Section D Cost Analysis, where we compare the computation and data requirement cost of REvolve with Eureka and RLHF on autonomous driving, as well as elaborate on the differences between REvolve and other existing approaches for autonomous driving; Section E Prompts, which presents system and operator (mutation, crossover) prompts for GPT-4 for REvolve and REvolve Auto; Section F Human Feedback Examples, which illustrates how human feedback allows GPT-4 to refine the reward function of individuals; Section G Additional Results, which lists the best reward functions generated from REvolve, and some additional results on RL convergence times and generalizability of the best reward functions.
@@
We also include our code and demonstrations of policies trained on best REvolve-designed reward functions in the supplementary material.
Appendix A Reinforcement Learning Preliminaries
A.1 Environments
A.1.1 Autonomous Driving
@@
Autonomous Driving experiments were conducted in the Neighborhood and Coastline environments of the AirSim simulator, a high-fidelity simulator developed for autonomous vehicle research (Shah et al., 2018). This urban simulation includes static features such as parked cars and dynamic obstacles like cars and moving animals, enabling a thorough evaluation of the RL agent in varied scenarios.
@@
Figure 5: AirSim simulator environment used for our autonomous driving experiments. The egocentric camera view displayed on the bottom left serves as the observations for the autonomous driving agent.
@@
Figure 6: AirSim generalization environments used for assessing the generalizability of REvolve-designed reward functions. The left and right pictures visualize Env 1 and Env 2, respectively.
Table 4: DDQN Hyperparameters used during training for all AirSim environments for the Autonomous Driving task.
@@
Training Algorithm. Due to the high dimensional observation space and convergence time, we adopted the Clipped Double Deep Q-Learning approach (Fujimoto et al., 2018), a refined version of Double Q-Learning (Van Hasselt et al., 2016). This method uses two separate neural networks to reduce overestimation bias by calculating the minimum of the two estimated Q-values, thus providing more reliable and stable training outcomes.
In addition to Clipped Double Deep Q-Learning, our framework integrates the Dueling Architecture (Wang et al., 2016), which improves the estimation of state values by decomposing the Q-function into two streams: one that estimates the state value and another that calculates the advantages of each action. Such structure allows the model to learn which states are valuable without the need to learn each action effect for each state.
Furthermore, we utilize Prioritized Experience Replay (PER) (Schaul et al., 2015) to focus on significant learning experiences. PER prioritizes transitions with higher Temporal Difference errors for replay, which accelerates learning and increases the efficiency of the memory used during training. Table 4 lists the RL model hyperparameters. Finally, inspired by foundational research, we have adapted our architecture to incorporate vision-based methods and the Dueling Architecture from Mnih et al. (2013) and Wang et al. (2016), enhancing temporal dynamics and action-value precision.
Figure 7: Humanoid Locomotion: A bipedal locomotion task where a simulated humanoid robot must maintain balance and walk as fast as possible (along the x-axis).
Figure 8: Adroit Hand: A complex robotic manipulation task requiring the dexterous control of a five-fingered robotic hand to rotate the latch and pull the handle to open a door.
A.1.2 Humanoid Locomotion
The Humanoid Locomotion environment is a complex bipedal locomotion task within the MuJoCo simulator (Todorov et al., 2012).
Observations. The observation space for the humanoid environment includes joint positions, joint velocities, and other information such as the center of mass and contact forces. The data is crucial for controlling balance and forward locomotion. The observation shape is a $376$D vector, representing the state of the humanoid at each timestep (Tassa et al., 2012).
Action Space. The action space is a $17$D continuous vector representing the torques applied to the humanoid’s joints, including those in the abdomen, hips, knees, shoulders, and elbows. This allows finer control of walking or running movements.
Objective. The goal is to walk or run forward without falling.
Termination. The episode ends either when the humanoid falls or $1000$ consecutive time steps have passed without falling (success).
Training Algorithm. We employed the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018), utilizing the implementation provided by the Stable Baselines3 (SB3) library (Raffin et al., 2021). SAC is an off-policy actor-critic method that optimizes a stochastic policy in an entropy-regularized reinforcement learning framework. By maximizing a trade-off between expected return and policy entropy, SAC encourages exploration and prevents premature convergence to suboptimal policies, making it well-suited for high-dimensional continuous action spaces characteristic of MuJoCo tasks.
We leveraged the default hyperparameters provided by SB3 for SAC, which have been empirically validated across various environments. Key hyperparameters included a learning rate of $3\times 10^{-4}$, a batch size of $256$, a discount factor ($\gamma$) of $0.99$, and an entropy coefficient ($\alpha$) automatically adjusted during training. The policy network employed the MlpPolicy architecture from SB3, consisting of multilayer perceptrons designed to process the high-dimensional state observations in MuJoCo environments effectively.
A.1.3 Adroit Hand
The Adroit Hand environment is another MuJoCo simulation task, which concerns dexterous robotic manipulation (Rajeswaran et al., 2017).
Observations. The observation space in the adroit hand environment consists of a $97$D vector. It includes joint positions for the hand, latch position, door position (hinge rotation), palm position, handle position, and the relative position between the palm and the handle. Additionally, the space contains joint velocities and forces acting on the hand. The binary door open indicator is also included to signify whether the door is fully opened or closed.
Action Space. The action space is a $30$D continuous vector that controls the absolute angular positions of the joints in the hand and arm. This enables dexterous manipulation tasks like unlatching and swinging open a door.
Objective. The agent’s goal is to manipulate the door latch and open the door, overcoming friction and bias torque forces.
Termination. The episode ends when the door is successfully opened or $400$ consecutive time steps have passed without opening (fail).
Training Algorithm. Same as humanoid locomotion.
Appendix B Fitness Functions
In this section, we provide details of (1) REvolve fitness function based on human evaluations, and (2) the automatic feedback (Auto) approach proposed in Eureka.